OpenStudy (anonymous):

Find the domain of the composite function f o g . Please show all of your work. f(x)=x+4; g(x)= 1/x+8

OpenStudy (rulnick):

If that's g(x) = 1/ (x+8) then the domain is all real numbers except x=-8.

OpenStudy (anonymous):

f(x)=x+4; g(x)= 1/x+8 -> f o g = (1/x+8 ) + 4 => Domain fog is R \ { -8 }